In five years, my daughter's age will be the same as twice her age 10 years ago. How old is my daughter now?
Answer by stanbon(75887)

In five years, my daughter's age will be the same as twice her age 10 years ago. How old is my daughter now?
---------
d+5 = 2(d-10)
----
d+5 = 2d - 20
d = 5+20
daughter's age = 25
--------------------
Cheers,
Stan H.
------------